# Nonreciprocal quantum phase transition in a spinning microwave magnonic system

### v1

@@ IV Conclusions
In summary, we have theoretically studied the role of rotation in engineering an nonreciprocal quantum phase transition in the cavity magnonic system consisting of a YIG sphere coupled to a spinning microwave resonator. By rotating the microwave resonator, we show that the introduced Sagnac effect can significantly modify the critical driving strengths for both second- and first-order quantum phase transitions and lead to a largely adjustable range of the critical driving strength. We further find that the critical driving strength of the phase transition relies on both the driving direction (or the rotation direction) and the spinning speed. In consequence, a highly tunable nonreciprocal quantum phase transition can be achieved based on the Sagnac effect. Our proposal provides an alternative route to achieve nonreciprocal quantum phase in microwave magnonical system and may find promising applications in designing nonreciprocal magnonic devices.

### v2

@@ IV Discussions and conclusions
In summary, we have theoretically studied the nonreciprocal QPT in the cavity magnonic system consisting of a YIG sphere placed in a spinning microwave resonator. By rotating the microwave resonator, we show that the introduced Sagnac effect can significantly modify the critical driving strengths for both second- and first-order QPTs and lead to a largely adjustable range of the critical driving strength. We further find that the critical driving strength of the phase transition relies on both the driving direction (or the rotation direction) and the spinning speed. In consequence, a highly tunable nonreciprocal QPT can be achieved based on the Sagnac effect. Our proposal provides an alternative route to achieve nonreciprocal quantum phase in microwave magnonical system and may find promising applications in designing nonreciprocal magnonic devices.